# VisionPAD: A Vision-Centric Pre-training Paradigm for Autonomous Driving

### v1

@@ 1 Introduction
Figure 1: Comparison with existing methods. (a) UniPAD employs volume rendering to reconstruct the multi-view depth maps and images of the current frame, using explicit depth maps for supervision. (b) In contrast, our proposed VisionPAD leverages only multi-frame, multi-view images for supervision, effectively learning motion and geometric representations through voxel velocity estimation and photometric consistency loss.
@@ 2 Related Works
Pre-training has been extensively explored in autonomous driving to improve scalability and adaptability across diverse driving environments. Existing approaches can be broadly categorized into supervised, contrastive, masked signal modeling, and rendering-based methods. Supervised approaches [31, 38, 36, 41] leverage annotated driving data to learn structured representations tailored for specific tasks. Contrastive approaches [19, 29, 32, 47] utilize positive and negative pairwise data to learn discriminative, view-invariant representations robust to variations within the driving scene. Masked signal modeling approaches [27, 28, 1, 16] reconstruct occluded or masked sensory signals, fostering a holistic understanding of scene semantics.
@@ 3.2 Volume Construction
Given $M$ historical multi-view images $\mathbf{I}=\{I_{i}\}^{M}_{i=1}$ as input to a representative vision-centric framework, VisionPAD employs a shared image backbone to extract 2D image features, resulting in a voxel feature representation $\mathbf{F}_{I}\in\mathbb{R}^{N\times H\times W\times C}$, where $N$, $H$ and $W$ denote the number of views, height and width of the images, respectively. Then, a view transformation [12, 20] lifts these 2D features into the 3D ego-centric coordinate system, generating volumetric features. Finally, a projection layer comprising several convolutional layers further refines the volumetric representation, which is denoted as $\mathbf{V}\in\mathbb{R}^{X\times Y\times Y\times Z\times C}$.
@@ 3.4 Self-supervised Voxel Velocity Estimation
Velocity Guidede Voxel Warping. As depicted in Fig. 3, an auxiliary velocity head, appended to the voxel features, regresses per-voxel velocity vectors in the world coordinates. After that, we approximate the per voxel flow by scaling the predicted velocities by the inter-frame time interval, enabling the transformation of voxel features from the current frame to an adjacent frame. Subsequently, we warp the voxel features to their estimated positions in the adjacent frame. The above process can be easily implemented through GridSample operator.
@@
UVTR+UniPAD CVPR’24 C        44.8 ↓0.2        38.5 ↑1.3
@@
UVTR+UniPAD CVPR’24 C        48.6 ↓0.2        40.5 ↑0.7
@@
Table 1: 3D object detection performance on the nuScenes val set. We benchmark against state-of-the-art methods across various modalities without test-time augmentation. Our approach achieves superior performance among existing vision-centric methods. “CS” denotes models trained with camera sweeps (two historical frames) as input. “CBGS” refers to the class-balanced grouping and sampling [52].
@@ 4.1 Experimental Settings
Implementation Details. Our implementation is based on MMDetection3D [7]. Following the same configuration as UniPAD [45], we use a ConvNeXt-small [24] as the default image encoder with an input resolution of $1600\times 900$ pixels. The constructed volume has shape $180\times 180\times 5$ within a valid perception range of [-54m, 54m] for the $X$ and $Y$ axis, and [-5m, 3m] for the $Z$ axis. The channel number is set to 256 for volume features. Data augmentation during pre-training includes random scaling and rotation, along with partial input masking (size=32, ratio=0.3 for images). The pre-training decoder consists of a 2-layer MLP and a 3D Convolution-BatchNorm-ReLU refine network for the velocity head, another 2-layer MLP with ReLU as an activation layer between them for predicting Gaussian primitives. The model is pre-trained for $12$ epochs using the AdamW optimizer, with initial learning rates set to $2\times 10^{-4}$, total batch size is 4. Fine-tuning follows the official downstream model configurations.
@@ 4.3 Comprehensive Analysis
Data Efficiency. A crucial benefit of pre-training is its capacity to enhance data efficiency in downstream tasks with limited labeled data. To further validate the benefits of our pre-training approach when pre-training data is abundant but downstream data is scarce, we conduct experiments using varying proportions of the nuScenes training set (10%, 25%, 50%, and 100%) to fine-tune a model pre-trained on the full nuScenes training dataset.

### v2

# 1FNii, Shenzhen   2SSE, CUHK-Shenzhen 3HKUST   4Huawei Noah’s Ark Lab

@@ 1 Introduction
Figure 1: Comparison with existing methods. (a) UniPAD employs volume rendering to reconstruct the multi-view depth maps and images of the current frame, using explicit depth maps for supervision. (b) In contrast, our proposed VisionPAD leverages only multi-frame, multi-view images for supervision, effectively learning motion and geometric representations through voxel velocity estimation and photometric consistency loss. re-pro. and sup. denote the re-projection and supervision, respectively.
@@ 2 Related Works
Pre-training has been extensively explored in autonomous driving to improve scalability and adaptability across diverse driving environments. Existing approaches can be broadly categorized into supervised, contrastive, masked signal modeling, and rendering-based methods. Supervised approaches [32, 39, 37, 42] leverage annotated driving data to learn structured representations tailored for specific tasks. Contrastive approaches [20, 30, 33, 48] utilize positive and negative pairwise data to learn discriminative, view-invariant representations robust to variations within the driving scene. Masked signal modeling approaches [28, 29, 1, 17, 15] reconstruct occluded or masked sensory signals, fostering a holistic understanding of scene semantics.
@@ 3.2 Volume Construction
Given $M$ historical multi-view images $\mathbf{I}=\{I_{i}\}^{M}_{i=1}$ as input to a representative vision-centric framework, VisionPAD employs a shared image backbone to extract 2D image features, resulting in a 2D feature representation $\mathbf{F}_{I}\in\mathbb{R}^{N\times H\times W\times C}$, where $N$, $H$ and $W$ denote the number of views, height and width of the images, respectively. Then, a view transformation [12, 21] lifts these 2D features into the 3D ego-centric coordinate system, generating volumetric features. Finally, a projection layer comprising several convolutional layers further refines the volumetric representation, which is denoted as $\mathbf{V}\in\mathbb{R}^{X\times Y\times Z\times C}$.
@@ 3.4 Self-supervised Voxel Velocity Estimation
Velocity Guided Voxel Warping. As depicted in Fig. 3, an auxiliary velocity head, appended to the voxel features, regresses per-voxel velocity vectors in the world coordinates. After that, we approximate the per voxel flow by scaling the predicted velocities by the inter-frame time interval, enabling the transformation of voxel features from the current frame to an adjacent frame. Subsequently, we warp the voxel features to their estimated positions in the adjacent frame. The above process can be easily implemented through GridSample operator.
@@
UVTR+UniPAD [46] CVPR’24 C        44.8 ↓0.2        38.5 ↑1.3
@@
UVTR+UniPAD [46] CVPR’24 C        48.6 ↓0.2        40.5 ↑0.7
@@
Table 1: 3D object detection performance on the nuScenes val set. Our approach achieves superior performance among existing vision-centric methods. “CS” denotes camera sweeps (two historical frames) input. “CBGS” refers to the class-balanced grouping and sampling [53]. $\dagger$ indicates our implementation. Pre-train Modal denotes the supervision modalities used during the pre-training stage.
@@ 4.1 Experimental Settings
Implementation Details. Our implementation is based on MMDetection3D [7]. Following the same configuration as UniPAD [46], we use a ConvNeXt-small [25] as the default image encoder with an input resolution of $1600\times 900$ pixels. The constructed volume has shape $180\times 180\times 5$ within a valid perception range of [-54m, 54m] for the $X$ and $Y$ axis, and [-5m, 3m] for the $Z$ axis. The channel number is set to 256 for volume features. Data augmentation during pre-training includes random scaling and rotation, along with partial input masking (size=32, ratio=0.3 for images). The pre-training decoder consists of a 2-layer MLP and a 3D Convolution-BatchNorm-ReLU refine network for the velocity head, another 2-layer MLP with ReLU as an activation layer between them for predicting Gaussian primitives. The model is pre-trained for $12$ epochs using the AdamW optimizer, with initial learning rates set to $2\times 10^{-4}$, total batch size is 4. Fine-tuning follows the official downstream model configurations without modification.
@@ 4.3 Comprehensive Analysis
Data Efficiency. A crucial benefit of pre-training is its capacity to enhance data efficiency in downstream tasks with limited labeled data. To further validate the benefits of our pre-training approach when pre-training data is abundant but downstream data is scarce, we conduct experiments using varying proportions of the nuScenes training set (10%, 25%, 50%, and 100%) to fine-tune a model pre-trained on the full nuScenes training dataset within the same epoches.
@@
Additionally, we found that using 2/3/4 anchors per voxel does not improve performance, with NDS dropping by 0.1/0.5/0.4.
@@ 5 Conclusion
Acknowledgements
This work was supported by Shenzhen General Program No. JCYJ20220530143600001, by the Basic Research Project No. HZQB-KCZYZ-2021067 of Hetao Shenzhen HK S&T Cooperation Zone, by NSFC with Grant No. 62293482, by Shenzhen-Hong Kong Joint Funding No. SGDX20211123112401002, by the Shenzhen Outstanding Talents Training Fund 202002, by Guangdong Research Project No. 2017ZT07X152 and No. 2019CX01X104, by the Guangdong Provincial Key Laboratory of Future Networks of Intelligence (Grant No. 2022B1212010001), by the Guangdong Provincial Key Laboratory of Big Data Computing, CHUK-Shenzhen, by the NSFC 61931024&12326610, by the Key Area R&D Program of Guangdong Province with grant No. 2018B030338001, by the Shenzhen Key Laboratory of Big Data and Artificial Intelligence (Grant No. ZDSYS201707251409055), by Tencent & Huawei Open Fund, and by China Association for Science and Technology Youth Care Program.
Appendix
Appendix A Additional Implementation Details
In this section, we elaborate on the implementation details of our proposed approach.
A.1 Evaluation Metrics
3D Objection Detection. In accordance with the evaluation protocols established in previous works [21, 12, 46], we utilize the NuScenes Detection Score (NDS) and mean Average Precision (mAP) as the primary metrics for assessing 3D object detection performance. The mAP is computed using the center distance on the ground plane rather than the 3D Intersection over Union (IoU), to match the predicted results and ground truth. Furthermore, the nuScenes dataset incorporates five types of true positive metrics (TP metrics), including ATE, ASE, AOE, AVE, and AAE for measuring translation, scale, orientation, velocity, and attribute errors, respectively. These TP metrics are defined for each class, and their means are computed across classes to yield mATE, mASE, mAOE, mAVE, and mAAE. Based on these TP errors, we define the TP score as $\mathrm{TP}_{\mathrm{score}}=\mathrm{max}(1-\mathrm{TP}_{\mathrm{error}},0)$. Subsequently, the nuScenes detection score (NDS) is computed as:
$\mathrm{NDS}=\frac{1}{10}\left[5\mathrm{mAP}+\sum{\mathrm{TP}_{\mathrm{score}}}\right],$ (9)
to encapsulate all aspects of the nuScenes detection comprehensively.
Semantic Occupancy Prediction. In the context of 3D semantic occupancy prediction, we adhere to the evaluation methodology employed by existing methods [36, 22, 13], utilizing mean intersection-over-union (mIoU) across all classes and intersection-over-union (IoU) as the evaluation metrics:
$\mathrm{mIoU}$ $=\frac{1}{|C^{\prime}|}\sum_{i\in C^{\prime}}\frac{{TP}_{i}}{{TP}_{i}+{FP}_{i}+{FN}_{i}},$ (10)
where $C^{\prime}=17$, $c_{0}$, $TP$, $FP$, $FN$ denote the number of semantic classes in the Occ3D-nuScenes dataset [36], the empty class, the number of true positive, false positive and false negative predictions, respectively.
Algorithm 1 The pseudocode of voxel velocity estimation.
0:  $\mathcal{V}_{t}$, $M$
0:  $\mathcal{V}_{t+n}$
  # predict the absolute flow for each voxel
  $\mathcal{F}_{t}\leftarrow\textbf{flow\_decoder}(\mathcal{V}_{t})$
  # transform the flow into absolute grid displacement to the future
  ^$\hat{\mathcal{F}}_{t+n}\leftarrow\Delta{t}\cdot n\cdot\mathcal{F}_{f}$
  # transform the absolute displacement to the future ego coordinate by using the current to future transformation
[…]
  # warping the current frame volume feature to the future
[…]
  return  $\mathcal{V}_{t+n}$
A.2 Details of Voxel Velocity Estimation Module
In the main paper, we have presented the effectiveness of the proposed self-supervised voxel velocity estimation module in capturing inherent temporal consistencies. Here we elaborate on more details of this module. The pseudocode for the voxel velocity estimation module is presented in Alg. 1. In this context, the inputs $\mathcal{V}_{t}$, $M$ represent the extracted volume features at the current frame and the pose transformation matrix from the current frame to a future frame, respectively. The output is the predicted volume features $\mathcal{V}_{t+n}$ in the future frame at time $n+t$. The process begins by predicting the absolute flow $\mathcal{F}_{t}$ for each voxel relative to the current frame coordinates, utilizing an MLPs-based flow decoder. Subsequently, the voxel flow is transformed into the voxel grid displacement for future frames using the formula $\Delta{t}\cdot n\cdot\mathcal{F}_{f}$, where $\Delta{t}$ and $n$ are the time interval between adjacent frames, and $n$ denotes the number of frames to be forecasted. In the nuScenes dataset, $\Delta{t}$ is set to $\frac{1}{12}$ by default, corresponding to the synchronized camera frame rate of 12 frames per second. The next step involves transforming the voxel displacement from the current frame to the future frame, relative to the future ego coordinate, by employing the pose transformation matrix. Finally, the future volume feature is obtained by warping the current volume feature with the predicted voxel displacement.
Primitives Variants Value NDS mAP
Mean Absolute - 26.8 25.9
Offset 0.25 27.3 26.5
 0.50 27.2 26.5
Scale Fixed 0.4 26.9 25.3
Learnable [0.1, 0.5] 27.3 26.5
[0.2, 0.8] 26.8 26.1
Rotation Fixed [1, 0, 0, 0] 27.2 26.7
Learnable - 27.3 26.5
Table 5: Ablation studies on various learning objectives for predicting Gaussian primitives.
Appendix B Additional Experimental Results
Due to the space constraint of the main paper, we provide more experimental results here to facilitate a more comprehensive understanding of our proposed method.
Effects of Gaussian Primitives. In VisionPAD, we present autonomous driving scenes as a set of Gaussian primitives $\{g_{k}=(\mu_{k},\Sigma_{k},\alpha_{k},c_{k})\}_{k=1}^{K}$, where each Gaussian $g_{k}$ is parameterized by its 3D position or mean $\mu_{k}\in\mathbb{R}^{3}$, a covariance $\Sigma_{k}$, an opacity $\alpha_{k}\in[0,1]$, and spherical harmonics (SH) coefficients $c_{k}\in\mathbb{R}^{k}$. The covariance $\Sigma_{k}$ is further formulated using a scaling matrix and rotation matrix to ensure its positive semi-definite. Therefore, it’s worth exploring various potential learning objectives for predicting these Gaussian primitives. To mitigate computational demands during training, this ablation study is confined to the 3D object detection task, employing $50\%$ of the training data during the pre-training phase and $25\%$ during the fine-tuning stage. Results are reported on the entire nuScenes validation set. The model used in this ablation study is consistent with the one employed in the main paper.
The results are presented in Tab. 5. We designed various experimental variants to explore the effects of different learning objectives. For each variant, we maintained the predictions of other primitives as consistent with those used in the main paper. We initially exploit the manners for the prediction of Gaussian mean parameters. By default, we transformed the Gaussian mean prediction problem into an offset prediction task using voxel centers as anchors. We also introduced an offset scaling factor to constrain the range of offset values. Consequently, a Gaussian mean can be obtained by:
$\mu_{k}=\mathbf{x}_{v}+(\mathrm{sigmoid}(\mathcal{O}_{k})-0.5)\cdot s,$ (11)
where $\mathbf{x}_{v}$ is the coordinate of an anchor point, $\mathcal{O}_{k}\in\mathbb{R}^{1\times 3}$ is the learnable offset, and $s$ is the scaling factor. We empirically set the $s$ to 0.25, as the same in our main paper, achieving a performance of 27.3 NDS and 26.5 mAP. When the scaling factor was increased to 0.5, no obvious performance changes were observed. Furthermore, instead of predicting position offset relative to the centers of volume anchors, we directly regressed the absolute $\mu_{k}$ via MLPs (second row in Tab. 5). This resulted in a performance decline in both NDS and mAP, from 27.3 NDS and 26.5 mAP to 26.8 NDS and 25.9 mAP. This indicates that it’s beneficial to predict Gaussian means by predicting the offsets relative to the anchor centers.
In relation to the Gaussian scale and rotation, our primary focus is on exploring the effects of their learnability. By default, the scale and rotation parameters are set to be learnable, with the scale constrained to the range of $[0.1,0.5]$. The final scale of Gaussian $k$ is calculated as:
$S_{k}=s_{l}+(s_{u}-s_{l})\cdot\mathrm{sigmoid}(\mathcal{S}_{k}),$ (12)
where $s_{l}$ and $s_{u}$ represent the lower and upper bounds of the scale, respectively, and $\mathcal{S}_{k}$ is the learnable scale parameter. We also experimented with a larger range of $[0.2,0.8]$, which resulted in a slight decrease in performance. When the scale and rotation is fixed by setting them as a constant of 0.4 (i.e. half of the voxel size) and $[1,0,0,0]$, i.e. the Gaussian degenerates into a spherical form. Fixing these parameters leads to varying degrees of performance decline, particularly notable when the scale is constant, resulting in a decrease in mAP from 26.5 to 25.3. This indicates that a fixed scale is inadequate for effectively modeling the scene.
Methods Decoder Memory Latency
UniPAD-C NeRF 1973MB 900ms
VisionPAD 3D-GS 134MB 70ms
Table 6: The speed analysis of our method with the UniPAD.
Figure 7: Visualization of 3D object detection and 3D semantic occupancy prediction results. Each detected object instance is depicted by a unique colored 3D bounding box. The legend at the bottom delineates the semantic classes of occupancy.
Speed Analysis. The NeRF-based methods often suffer from substantial GPU memory consumption and slow rendering speeds. Furthermore, the resource consumption in NeRF-based methods notably increases with larger rendering sizes. In contrast, our 3D-GS-based approach maintains faster rendering speeds irrespective of rendering size. As presented in Tab. 6, we compare the efficiency and memory consumption between UniPAD-C with ours when rendering the same $360\times 640$ images. We conduct this experiment in a single NVIDIA A100 GPU. Our approach distinctly results in an approximate 93.2% reduction in memory usage, alongside an approximate 92.2% decrease in rendering latency. This validates the potential of the 3D-GS-based pre-training paradigm.
UniPAD with V.V. Est. and P.C. To further validate the effectiveness of the proposed photometric consistency (P.C.) and the self-supervised voxel velocity estimation (V.V. Est.), we apply these two strategies in the UniPAD. The results are shown in Tab. 7, our proposed modules achieve significant improvements on UniPAD (+3.3%/+5.0%). It is worth noting that 3D-GS renders the entire image with lower computational cost. Compared to ray-sampling approach in UniPAD, 3D-GS increases the effective pixels for photometric consistency supervision, thereby achieving better results.
Metrics UniPAD w/ V.V. Est. and P.C. VisionPAD
NDS 22.3 25.6 27.3
mAP 18.3 23.3 26.5
Table 7: The effectiveness and generalizability of proposed proposed photometric consistency (P.C.) and the self-supervised voxel velocity estimation (V.V. Est.) strategies for pre-training.
Computation Cost of Gaussian Filtering. As illustrated in Tab. 8, we demonstrate that using the Gaussian filter can effectively reduce memory usage and improve speed.
Model Memory Latency NDS
w/o filter 186MB 90ms 26.8
VisionPAD 134MB 70ms 27.3
Table 8: The computation cost analysis of Gaussian filtering operation.
Appendix C Additional Visualization Results
As illustrated in Fig. 7, we visualize an additional scene to qualitatively assess the efficacy of our method. To facilitate an intuitive comparison of the perception results, we simultaneously demonstrate the 3D object detection and 3D semantic occupancy prediction within the same scene. The 3D object detection results are predicted by the UVTR [18], while the 3D semantic occupancy prediction results are obtained by the BEVDet [12]. Both of them are pre-trained by our method.
From the results, we can not only discern the desirable performance of both 3D object detection and occupancy prediction but also highlight the respective strengths and limitations inherent in these two representations. The dense occupancy representation, is capable of depicting arbitrary irregular objects and background obstacles in a detailed voxel format, but tends to demand higher computational resources. Conversely, 3D bounding boxes primarily represent foreground objects in a sparse format. In this scene, despite the accurate detections of most objects, some false positive bounding boxes are observed, and the orientation is also not accurate enough to describe the excavator at the front right. In comparison, the occupancy predictions offer a more satisfactory representation of these critical objects. Encouragingly, as the visualization results indicate, our pre-training paradigm enhances performance across both tasks. It further underscores the efficacy and adaptability of our proposed method.